# Reorganizing attention-space geometry with expressive attention

### v2

@@ 6 Discussion
Repeatedly, we observed that initial performance gains flatten out rapidly. This happens in particular for small model sizes, but also for DPA in cases where the available resources are sufficient for EA to solve the problem at hand exactly. We argue that the resulting stationary performance plateau indicates the presence of a local minimum in loss landscape. From this trap models may escape either by embedding the task at hand into into a larger dimensional parameter space, or by improved design. For all cases, the level of the observed stationary performance plateau was independent of initial conditions and training details. This led us to the conclusion that the associated local minimum in loss landscape is structural, viz that it corresponds to a heuristic strategy. To examine how this heuristic strategy works would an interesting research question, which is however beyond the scope of the present study.
@@
Inductive reasoning is one of the big challenges of large language models (?, ?). It is encouraging that one of the building blocks, the logical if-statement included, can be extracted and encoded when effective attention is used. Our analysis, which involved logical task switching within the NT-R framework (9), is however only a first indication and it remains to be seen to which extent EA may raise performance in this field.
Attention space geometry
A somewhat intriguing insight emerging from the present study concerns the nature of the attention mechanism. It is clearly possible that qualitative different realizations of the underlying principle may show somewhat comparative performance levels. What matters in first place is the core working principle, dynamic information routing, and only in second place how this principle is implemented in attention space, e.g. using dot-product or expressive attention.

### v3

@@ Discussion
Repeatedly we observed that initial performance gains flatten out rapidly. This happens in particular for small model sizes, but also for DPA in cases when the available resources are sufficient for EA to solve the problem at hand exactly. We argued that the resulting stationary performance plateau indicates the presence of a local minimum in loss landscape. From this trap models may escape either by an embedding into into a larger dimensional parameter space, or by an improved design. For all cases the level of the observed stationary performance plateau was independent of initial conditions and training details. This led us to the conclusion that the associated local minimum in loss landscape is structural, viz that it corresponds to a heuristic strategy. To examine how this heuristic strategy works would an interesting research question, which is however beyond the scope of the present study.
@@
Inductive reasoning is one of the big challenges of large language models (Li et al. 2024b; Su et al. 2023). It is encouraging that one of the building blocks, the logical if-statement included, can be extracted and encoded when effective attention is used. Our analysis, which involved logical task switching within the NT-R framework (9), is however only a first indication and it remains to be seen to which extent EA may raise performance in this field.